Annilee's present age is two-thirds of Jessie's present age. In 4 years the sum of their ages will be 73 years. Find their present ages.
Answer by Maths68(1474)

Let
Present age of Annilee=a
Present age of Jessie=j
Annilee's present age is two-thirds of Jessie's present age
a= (2/3)j
a=2j/3 ................(1)
In 4 years means we have to add 4 in the present ages of both
a+4=j+4
then
the sum of their ages will be 73 years
(a+4)+(j+4)=73
a+4+j+4=73
a+j+8=73
a+j=73-8
a+j=65..............(2)
Put the value of a from (1) to (2)
a+j=65
2j/3 + j = 65
Multiply by 3 both sides of above equation
2j+3j=195
5j=195
Divide by 5 both sides of the equations
5j/5=195/5
j=39
Put the value of j in (1)
a=2j/3
a=(2*39)/3
a=78/3
a=26
Present age of Annilee= 26
Present age of Jessie= 39